OpenStudy (anonymous):

Jerry is rolling a die and flipping a coin. What is the probability that he will roll an even number and the coin will land on tails? 1/5

OpenStudy (lgbasallote):

3/6 x 1/2 i guess

Parth (parthkohli):

When we use the word 'AND', then we must multiply the probabilities. P(Getting an even number) = 1/2 P(Getting a tail) = 1/2 \(\Large \color{MidnightBlue}{\Rightarrow ({1 \over 2})^2 }\)

OpenStudy (lgbasallote):

3/6 is the probability of rolling even 1/2 is the probability of tail

Parth (parthkohli):

@lgbasallote 3/6 simplified is 1/2

OpenStudy (lgbasallote):

i know...but for simplicity sake i used 3/6

Parth (parthkohli):

Simplicity comes when we actually simplify it.

OpenStudy (lgbasallote):

it's not my habit to explain something and jump steps

OpenStudy (lgbasallote):

i believe in thE KISS principle

Parth (parthkohli):

Wow.